OpenStudy (anonymous):

Solve the equation. 4/9 n + 6= 4/3 a) -4 2/3 b) -10 1/2 c) 10 1/2 d) 16 1/2 I need help.

OpenStudy (timothywhitaker):

add the 6 to the 4/3 then divide your answer by 4/9

OpenStudy (anonymous):

maybe you want to subtract first

OpenStudy (anonymous):

if it was me, i would multiply both sides by 9and start with \[4n+54=12\] and solve that in two steps

OpenStudy (anonymous):

the way my gradpoint showed me was that you had to do the reciprocal of 4/9n which would be 9/4. I have the equation up now as 9/4 (4/9n) + 6= 9/4 (4/3)

OpenStudy (timothywhitaker):

thats wat i ment i got B

OpenStudy (anonymous):

-20/2 which euqals -10 1/2

OpenStudy (anonymous):

*equals